OpenStudy (anonymous):

How many solutions are there to the equation below? 17x - 8 = 3x + 16 A. Infinitely many B. 0 C. 1

OpenStudy (unklerhaukus):

have you tried solving the equation for x?

OpenStudy (unklerhaukus):

\[17x - 8 = 3x + 16\]\[17x - 3x = 8 + 16\]